OpenStudy (anonymous):

Find the monthly payment for the loan. New car financing of 3.3% on a 30-month $12,850 loan. I know to use the amortization formula, but the 30-months totally threw me off and I'm not sure what number to use for 'n' and 't'. Please help!

jimthompson5910 (jim_thompson5910):

what is the amortization formula your book gives you?

OpenStudy (anonymous):

\[m=\frac{ P(\frac{ r }{ n }) }{ 1-(1+\frac{ r }{ n })^{-nt} }\]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

ok t is time in years 30 months = 30/12 = 5/2 = 2.5 years so t = 2.5 ----------------------------------- n is the payment frequency. In this case, it happens monthly which means n = 12 r = 3.3/100 = 0.033 since 3.3% = 0.033

OpenStudy (anonymous):

Thank you so much for clearing that up for me!!! It makes so much more sense now!

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking now

OpenStudy (anonymous):

@jim_thompson5910 Thank you! I got the correct answer! Do you mind helping me with another problem? How much interest (to the nearest dollar) would be saved on the following loan if the home were financed for 15 rather than 30 years? A $128,000 home bought with a 20% down payment and the balance financed for 30 years at 8.5%

jimthompson5910 (jim_thompson5910):

how much is paid for the down payment?

OpenStudy (anonymous):

I'm not sure that was all the information I was given

jimthompson5910 (jim_thompson5910):

well it says "20% down payment" on the $128,000 home

OpenStudy (anonymous):

I don't know how to figure that out

jimthompson5910 (jim_thompson5910):

you take 20% of the home's value

OpenStudy (anonymous):

$6400?

OpenStudy (anonymous):

or $25600?

jimthompson5910 (jim_thompson5910):

it should be 0.20*128,000 = 25,600

jimthompson5910 (jim_thompson5910):

you have to pay $25,600 up front before you can even think about getting the mortgage

jimthompson5910 (jim_thompson5910):

the remaining about you didn't pay, 128,000 - 25,600 = 102,400 is the amount you are financed. Ie, it is the amount that is lent to you.

jimthompson5910 (jim_thompson5910):

amount*

jimthompson5910 (jim_thompson5910):

so P = 102400

jimthompson5910 (jim_thompson5910):

8.5% = 0.085 so r = 0.085

jimthompson5910 (jim_thompson5910):

n = 12 because mortgages are paid monthly t = 30 (at first)

OpenStudy (anonymous):

is $787.37 the correct answer? And then I would do the same with t=15

jimthompson5910 (jim_thompson5910):

that is the monthly payment for the 30 yr mortgage

jimthompson5910 (jim_thompson5910):

how much total will you pay back over the entire 30 yr period?

OpenStudy (anonymous):

is that done with an annuity formula?

jimthompson5910 (jim_thompson5910):

nope, multiply the monthly payment by 360 why 360? because there are 360 months in 30 years 30 years = 30*12 = 360 months

jimthompson5910 (jim_thompson5910):

simple example: let's say you pay $100 a month for 30 months, you would pay back a total of 100*30 = 3,000 dollars

OpenStudy (anonymous):

$283,452.27

jimthompson5910 (jim_thompson5910):

I'm getting $283,453.20

OpenStudy (anonymous):

I redid it and got the same answer as you

jimthompson5910 (jim_thompson5910):

now we subtract the initial amount borrowed ($102,400) from the total ($283,453.20)

jimthompson5910 (jim_thompson5910):

so... 283,453.20 - 102,400 = 181,053.2

jimthompson5910 (jim_thompson5910):

this means that the total amount of interest and interest only is $181,053.20

jimthompson5910 (jim_thompson5910):

again this is all for the 30 yr mortgage

OpenStudy (anonymous):

Okay so that completes the 30 yr part? Do I do the same for the 15 yr?

jimthompson5910 (jim_thompson5910):

yes, follow the same steps for the 15 yr one

jimthompson5910 (jim_thompson5910):

then subtract the two interest only figures to get the amount you save

OpenStudy (anonymous):

And that will be the answer?

jimthompson5910 (jim_thompson5910):

correct. After you round to the nearest dollar

OpenStudy (anonymous):

$101946.60

jimthompson5910 (jim_thompson5910):

I'm getting the same answer. Nice work.

OpenStudy (anonymous):

Thank you so much for your help!! I really appreciate it!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

@jim_thompson5910 I really hate to be a bother, but if you don't mind could you help me with two more problems? I tired to do both of them, but I just can't seem to figure them out! 1. Melissa agrees to contribute $500 to the alumni fund at the end of each year for the next 5 years. Shannon wants to match Melissa's gift, but he wants to make a lump-sum contribution. If the current interest rate is 7.5% compounded annually, how much should Shannon contribute to equal Melissa's gift? (Round your answer to the nearest cent.) 2. A $1,000,000 lottery prize pays $50,000 per year for the next 20 years. If the current rate of return is 8.25%, what is the present value of this prize? (Assume the lottery pays out as an ordinary annuity. Round your answer to the nearest cent.)

jimthompson5910 (jim_thompson5910):

ok let me think them over for a moment

jimthompson5910 (jim_thompson5910):

I think for the first one, you use the future value of annuity formula. Do you have that formula with you?

OpenStudy (anonymous):

Yeah I do

jimthompson5910 (jim_thompson5910):

what formula do they give you?

OpenStudy (anonymous):

\[P= m(\frac{ 1-(1+\frac{ r }{ n }^{-nt}) }{\frac{ r }{ n } })\]

jimthompson5910 (jim_thompson5910):

m = 500 is the amount Melissa pays per year r = 0.075 (since 7.5% = 7.5/100 = 0.075) n = 1 (compounded annually) t = 5 (for five years) plug those values into the formula to get ???

OpenStudy (anonymous):

$2022.94

jimthompson5910 (jim_thompson5910):

I got that as well

OpenStudy (anonymous):

It showed up correct on my homework! :D

jimthompson5910 (jim_thompson5910):

that's the amount Shannon needs to contribute

jimthompson5910 (jim_thompson5910):

As for the next one, you'll most likely use the present value formula. What formula do they give you?

OpenStudy (anonymous):

The same one as before?

OpenStudy (anonymous):

Nevermind I found it, just a second

jimthompson5910 (jim_thompson5910):

no I don't think you'll use the future value formula

OpenStudy (anonymous):

\[P=A(1+\frac{ r }{ n })^{-nt}\]

jimthompson5910 (jim_thompson5910):

hmm does it give you a present value of annuity formula?

OpenStudy (anonymous):

The present value of annuity formula that I was given is the one I used from before

jimthompson5910 (jim_thompson5910):

so this? \[P= m(\frac{ 1-(1+\frac{ r }{ n }^{-nt}) }{\frac{ r }{ n } })\]

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

it seems odd how you'd use that formula though

OpenStudy (anonymous):

It's okay if you can't figure out how to solve it!

OpenStudy (anonymous):

Would it be the ordinary annuity formula?

jimthompson5910 (jim_thompson5910):

I did find this page http://www.financeformulas.net/Present_Value_of_Annuity.html

jimthompson5910 (jim_thompson5910):

I wonder if you are given a similar formula?

OpenStudy (anonymous):

I don't have a formula like that, but maybe it will work

jimthompson5910 (jim_thompson5910):

using that formula, the inputted values are P = 50,000 r = 0.0825 t = 20

OpenStudy (anonymous):

would n be 12?

jimthompson5910 (jim_thompson5910):

well I'm using the formula given on this page http://www.financeformulas.net/Present_Value_of_Annuity.html oh I completely glossed over n for some reason. The variable n represents the number of periods. In this case, there are 20 periods (20 years). One period is one payment. So in a way, this page is treating "t" like "n". A bit confusing if you ask me.

jimthompson5910 (jim_thompson5910):

n = 20 in this case because 20 payments are made

jimthompson5910 (jim_thompson5910):

each payment is coming from the lottery commission and it goes to the lottery winner

OpenStudy (anonymous):

$481907.39

jimthompson5910 (jim_thompson5910):

got the same

OpenStudy (anonymous):

Yay!!! THANK YOU SOOOOOOO MUCH!! I cannot thank you enough, I would have no idea what to do without you!

jimthompson5910 (jim_thompson5910):

I'm glad I could help out and that it is making sense now.